OpenStudy (samsan9):

Let csc t= sqrt (5) with t in Q2 and find the following. cos2A and Sec2A

OpenStudy (solomonzelman):



OpenStudy (samsan9):

which formula do we use for both?

OpenStudy (samsan9):

the missing side is 2

OpenStudy (solomonzelman):

Then, you have \(\color{black}{ {\rm adj=\sqrt{hyp^2-opp^2}=\sqrt{(\sqrt{5})^2-1}=\sqrt{4}=2}}\). Yes, you are correct.

OpenStudy (solomonzelman):

Then, you know that \(\color{black}{ \cos(t)={\rm adj/hyp}}\) \(\color{black}{ \sec(t)={\rm hyp/adj}}\)

OpenStudy (solomonzelman):

The adjacent side is -2.

OpenStudy (samsan9):

Wait so I know that Cos2A is =cos^2A-sin^2A but Sec is the hardest one for me

OpenStudy (solomonzelman):

(Due to the angle being in the second quadrant)

OpenStudy (solomonzelman):

well, sec = 1 /cos

OpenStudy (solomonzelman):

So, you got: \(\color{black}{ \cos(2t)=2\cos^2(t)-1=2\left(\frac{-2}{\sqrt{5}}\right)^2-1 }\).

OpenStudy (solomonzelman):

and you know that \(\color{black}{ \displaystyle \sec(2t)=\frac{1}{\cos(2t)}=\frac{1}{2\cos^2(t)-1}=\frac{1}{2\left(\frac{-2}{\sqrt{5}}\right)^2-1 } }\).

OpenStudy (samsan9):

thank you

OpenStudy (solomonzelman):

\({\rm yw}{\tiny~}:\)\()\)